OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (anonymous):

Wouldn't it be C.

OpenStudy (anonymous):

y-y1=m(x-x1) (x1,y1)=(-2,-3)m=-6

OpenStudy (anonymous):

acc to me it will be A ... give out u r working

OpenStudy (anonymous):

y- -3=6(x- -2) y- -3= 6x - -12 +3 +3 y= 6x + 15

OpenStudy (anonymous):

It was subtraction not addition

OpenStudy (anonymous):

y- (-3)=-6(x- (-2)) y+3=- 6x -12 -3 -3 y= -6x - 15 got it

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

u welcome ^_^

OpenStudy (anonymous):

Can you do one more please